As a current MBA student, your friend has thought it best to contact you for some retirement advice. He would like to retire in exactly 40 years. Your friend has the following retirement goals: Once he has retired, he would like to be able to withdraw R1,000,000 per year for 25 years in order to live comfortably (assume withdrawals are made at the end of each period). He would also like to leave an inheritance of R2,000,000 for his family at the end of the 25 years. Assume that the interest rate that he will be able to receive on the investment during his retirement is 10%

Question: Calculate the amount of money that your friend will require at his retirement date. Show all your
Answer #1
The amount of money required on retirement date is the sum
of the PV of the R2,000,000 that has to be left at the end of 25
years after retirement and the PV of the R1,000,000 to be
withdrawn at the end of each of the 25 years after retirement.
PV of R2,000,000 = 2000000/1.1^25 = $       1,84,592
PV of the annual withdrawal [annuity] = 1000000*(1.1^25-1)/(0.1*1.1^25) = $     90,77,040
Amount of money required at retirement date $     92,61,632